Find the potential º when F =-V¢, a conservative force of field defined by F = (3x*yz - 3yli +(rºz – 3x); +(x*y+22)
Answer #1

Given F=-06 -10=<Dy, 0) pis potential function F = conservative force Given E= (3x²yz - 3y) ↑ + (x32-34) ↑ + cxly +22) x = 3x